# Training Long-Context, Multi-Turn Software Engineering Agents with Reinforcement Learning

### v1

@@ Abstract
Research on applications of Reinforcement Learning (RL) to Large Language Models (LLMs) has mostly been focused on single-turn problems, such as mathematical reasoning or single-shot code generation. While these problems can be viewed as token-level multi-turn MDPs, this view corresponds to a degenerate case of multi-turn interaction where the environment provides no feedback. This contrasts with many real-world domains, such as software engineering (SWE), which require rich multi-turn interactions with a stateful environment that responds to each action with a non-trivial observation.
To bridge this gap, we demonstrate the successful application of RL to this general regime. Using a modified Decoupled Advantage Policy Optimization (DAPO) algorithm, we train an agent based on Qwen2.5-72B-Instruct to solve real-world software engineering tasks. Our approach increases the agent’s success rate on the SWE-bench Verified benchmark from a 20% rejection fine-tuned baseline to 39%, without relying on any teacher models. On SWE-rebench, our agent matches or outperforms leading open-weight models such as DeepSeek-V3-0324 and Qwen3-235B-A22B using an identical scaffolding, offering a viable path toward building more capable autonomous agents for complex real-world problems based on open models.
@@ 1 Introduction
Figure 1: Performance of the RL-trained agent over training iterations. To accelerate evaluations, Pass@1/10 are computed on a random subset of 50 tasks from SWE-bench Verified and averaged over 10 independent runs to reduce variance. Pass@1 is also reported together with the standard error of the mean.
Large Language Models are increasingly deployed within autonomous agents in complex, real-world domains. Software engineering is an especially compelling application area, promising substantial economic impact through automation of debugging, code generation, and software maintenance tasks. However, current approaches to developing effective SWE agents rely predominantly on one of three strategies: (i) combining sophisticated scaffolding with proprietary LLMs (AugmentCode 2025; Refact AI 2025; Trae 2025), (ii) leveraging extensive inference-time scaling techniques (Antoniades et al. 2025), or (iii) supervised fine-tuning (SFT) of open-weight models using demonstrations from stronger teacher models (Yang et al. 2025; Zeng et al. 2025; Pan et al. 2025). While these methods have yielded strong initial results, they are often resource-intensive and depend on powerful, proprietary models. This reality underscores the need for methods that can build comparably effective systems from smaller, open-weight models. Reinforcement learning offers a promising alternative by directly optimizing an agent’s policy through interaction with a responsive environment, potentially achieving stronger performance without reliance on teacher models.
@@
Figure 2: Illustration of task structure differences between bandit-style problems (top, e.g., math) and POMDPs (bottom, e.g., software engineering). In bandit settings, the agent takes a single action to produce a final solution based on an initial observation. In contrast, POMDPs require a multi-step interaction loop where the agent repeatedly takes actions and interprets new environmental feedback to guide its subsequent decisions.
@@
  • Data scalability and fidelity: Generating high-quality trajectories requires the reproduction of specific repository states in controlled environments, which limits dataset scale. Large-scale datasets such as SWE-smith (Yang et al. 2025) and SWE-rebench (Badertdinov et al. 2025) begin to address this gap; we primarily build on the latter.
@@
In this paper, we address these challenges by developing a complete RL pipeline tailored explicitly to interactive SWE tasks. Our core contributions include:
  • A scalable RL framework, based on a modified DAPO algorithm (Yu et al. 2025), specifically adapted to handle the demands of long-horizon, multi-turn SWE scenarios.
  • Empirical demonstration of RL effectiveness by training a Qwen2.5-72B-Instruct agent that achieves approximately 39% success on the SWE-bench Verified benchmark, doubling the baseline performance of a rejection-fine-tuned agent. Moreover, our agent matches or surpasses top open-weight models, such as DeepSeek-V3-0324 and Qwen3-235B-A22B, on the SWE-rebench benchmark (Table 1).
  • Detailed analysis of our RL training methodology, including algorithmic modifications, hyperparameter settings, key findings and discussion of promising future directions for applying RL to LLM-based agents in interactive, stateful environments.
Our results show that RL can be successfully applied to general, interactive environments, advancing the capabilities of open-weight LLM agents beyond single-turn benchmarks and toward real-world scenarios, where future autonomous agents are expected to operate.
@@ 2.1 Task Formulation
We formalize the task of autonomous SWE agent as a Partially Observable Markov Decision Process (POMDP) (Murphy 2025), defined by the tuple $\langle\mathcal{Z},\mathcal{A},\Omega,T,\mathcal{O},R,\gamma\rangle$:
@@
Our goal is to find a policy $\pi_{\theta}$ that maximizes the expected cumulative reward. The cumulative reward for a trajectory $\tau$ is the sum of rewards over all its steps, $G(\tau)=\sum_{t=0}^{|\tau|-1}R(z_{t},a_{t})$, where $|\tau|$ denotes the number of actions in the trajectory. Given our sparse reward structure where immediate rewards $R(z_{t},a_{t})$ are zero for all non-terminal steps, this simplifies to the terminal reward, which we denote as $R(\tau)$. $R(\tau)$ is $1$ if the final proposed patch passes the validation test suite, and 0 otherwise. We use this terminal reward as the sole signal for policy optimization.
@@ 2.2 From PPO to DAPO
Reinforcement learning has traditionally been dominated by Proximal Policy Optimization (PPO) (Schulman et al. 2017), which employs a learned critic to estimate the advantage of each generated action. While powerful, PPO introduces extra overhead and sensitivity to hyperparameter tuning. In our setting, the PPO objective can be the following:
$\mathcal{J}_{\text{PPO}}(\theta)=\mathbb{E}_{\begin{subarray}{c}o_{0}\sim\mathcal{D},\\ \tau\sim\pi_{old}(\cdot|o_{0})\end{subarray}}\Biggl{[}\frac{1}{|\tau|}\sum_{t=0}^{|\tau|-1}\frac{1}{|a_{t}|}\sum_{k=1}^{|a_{t}|}\min\Bigl{(}\\ \rho_{t,k}(\theta)A_{t,k},\text{clip}(\rho_{t,k}(\theta),1-\varepsilon,1+\varepsilon)A_{t,k}\Bigr{)}\Biggr{]},$ (2)
where $o_{0}\sim\mathcal{D}$ indicates that initial observations are sampled from the distribution of tasks in our dataset. $\rho_{t,k}(\theta)\triangleq\frac{\pi_{\theta}(a_{t,k}|h_{t},a_{t,<k})}{\pi_{\theta_{\text{old}}}(a_{t,k}|h_{t},a_{t,<k})}$ is the probability ratio for the $k$-th token in action $a_{t}$, and $A_{t,k}$ is its corresponding advantage estimate.
Group-Relative Policy Optimization (GRPO) (Shao et al. 2024) eliminates the need for a learned advantage estimator. Instead, it computes a Monte Carlo estimate of the advantage for each trajectory by normalizing its terminal reward against the mean and standard deviation of rewards from a group of trajectories sampled from the same policy.
@@
where $\sigma_{R}$ is the standard deviation of rewards in the group and $\delta$ is a small constant for numerical stability. This single advantage value $\hat{A}^{(i)}$ is then applied to all tokens within the trajectory $\tau^{(i)}$.
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{\begin{subarray}{c}o_{0}\sim\mathcal{D},\\ \{\tau^{(i)}\}_{i=1}^{G}\sim\pi_{old}(\cdot|o_{0})\end{subarray}}\Biggl{[}\\ \frac{1}{G}\sum_{i=1}^{G}\frac{1}{|\tau^{(i)}|}\sum_{t=0}^{|\tau^{(i)}|-1}\frac{1}{|a_{t}^{(i)}|}\sum_{k=1}^{|a_{t}^{(i)}|}\min\Bigl{(}\\ \quad\rho_{t,k}^{(i)}(\theta)\hat{A}^{(i)},\text{clip}(\rho_{t,k}^{(i)}(\theta),1-\varepsilon,1+\varepsilon)\hat{A}^{(i)}\Bigr{)}\Biggr{]}.$ (4)
@@
  • Dynamic sampling: Sampled commands that carry no learning signal (i.e., $\hat{A}^{(i)}=0$) are filtered out, focusing computation on effective updates.
  • Soft overlong punishment: Specifically, when the response exceeds a predefined length threshold, a gradual penalty within a specified interval is applied. Within this interval, the longer the response, the greater the punishment it receives. This penalty is added to the original test-based reward, thereby signaling to the model to avoid excessively long responses.
@@
$\mathcal{J}_{\text{DAPO}}(\theta)=\mathbb{E}_{\begin{subarray}{c}o_{0}\sim\mathcal{D},\\ \{\tau^{(i)}\}_{i=1}^{G}\sim\pi_{old}(\cdot|o_{0})\end{subarray}}\Biggl{[}\\ \frac{1}{\sum_{i=1}^{G}\sum_{t=0}^{|\tau^{(i)}|-1}|a_{t}^{(i)}|}\sum_{i=1}^{G}\sum_{t=0}^{|\tau^{(i)}|-1}\sum_{k=1}^{|a_{t}^{(i)}|}\min\Bigl{(}\\ \quad\rho_{t,k}^{(i)}(\theta)\hat{A}^{(i)},\text{clip}(\rho_{t,k}^{(i)}(\theta),1-\varepsilon_{low},1+\varepsilon_{high})\hat{A}^{(i)}\Bigr{)}\Biggr{]}.$ (5)
Our implementation builds on DAPO, but introduces some modifications for the multi-turn SWE setting, further discussed in detail.
@@ 2.3 Agent Scaffolding
The SWE agent in our work employs a ReAct-style loop (Yao et al. 2023), interacting with the environment through a predefined set of tools. The entire action–observation history conditions every decision. We provide an evaluation infrastructure supporting over 7,000 curated SWE tasks, each with a natural-language issue description and a reproducible environment.
Our SWE agent interacts with an environment using the following tools:
@@
Each SWE task includes a GitHub-style issue with a natural language description, a failing test suite that evaluates final patch correctness, and a sandboxed environment initialized from a repository snapshot.
3 Reinforcement Learning for Multi-Turn Agents
This section outlines our training methodology, consisting of carefully curated data and a two-phase training pipeline optimized explicitly for multi-turn SWE tasks.
@@ 3.1 Data
  • Deterministic tests: Remove tasks exhibiting flakiness upon repeated executions of tests (50 trials), ensuring stable training signals. Non-deterministic test behavior occurs mostly due to external service calls or floating-point inaccuracies.
For evaluation, we use the standard SWE-bench Verified benchmark, a 50-problem random subset (referred to as Verified-50) for faster intermediate evaluations, and the latest monthly splits of SWE-rebench (May-June) which are not included in the training set, ensuring a fair and decontamination-free comparison. The full list of Verified-50 problems can be found in Appendix B.
@@ 3.2 Phase 1: Rejection Fine-Tuning (RFT)
We start from the open-weight Qwen2.5-72B-Instruct model. Out of the box it reaches only $\sim$11% on SWE-bench Verified with the dominant failure being incorrect instruction following, resulting in commands that are not in the appropriate format.
To address this problem and warm up the model for the RL stage, we perform rejection fine-tuning (RFT). First, we run the initial checkpoint 10 times on the 7,249 SWE-rebench tasks and keep only those trajectories whose patches pass the test suite. This yields a set of 6,548 successful trajectories, on which we run a single epoch of supervised fine-tuning. During this epoch, we mask every agent turn that triggered an environment-formatting error, thereby focusing the loss on valid actions and improving adherence to the tool structure (see Figure 3). After RFT, the model’s accuracy rises to $\sim$20% (Table 1). All hyperparameters for this and subsequent stages are listed in the training details section.
Figure 3: An example trajectory from the agent’s interaction used in RFT. Only green (successful) assistant turns contribute to training loss.
@@ 3.3 Phase 2: Multi-Turn RL
The core of our work involves applying RL over thousands of problems in an iterative loop. Each RL iteration includes:
@@
  • Reward computation: The final reward signal, $R_{\text{final}}(\tau)$, combines the binary success reward from testing with a penalty for trajectory length. We add a linear penalty for exceeding a certain number of steps. The reward is computed as follows:
@@
    Here, $R(\tau)\in\{0,1\}$ is the terminal reward defined in the previous section, $L_{\text{thr}}$ and $T_{\text{max}}$ are hyperparameters representing the penalty threshold and maximum episode length, respectively.
@@
  • Optimization: We update model’s all parameters using DAPO’s clipped token-level objective.
We begin RL training at a 65k context length, already beyond the default 32k of most open-weight LLMs, but still insufficient for repositories with long stack traces or diff histories. Once performance plateaus at approximately 32%, we switch to a longer context setting where we increase the context window to 131k and double the maximum number of agent turns $T_{\text{max}}$ from 40 to 80. For stable updates in this second stage, we also decrease the high clip bound $\epsilon_{\text{high}}$, increase the batch size, and lower the number of instances sampled on each iteration, i.e., shifting training closer to on-policy mode (see Table 3).
As part of the transition between stages, we perform a one-time curriculum adjustment by refreshing the training pool. We define a task’s cumulative resolution rate as its success rate over all prior training iterations. Tasks are removed from the pool if their cumulative resolution rate exceeds $\frac{2}{3}$ (indicating the task can be solved reliably) or remains at zero (indicating it is likely unsolvable). This filtering reduces the training pool from 7,249 to 2,028 tasks. This strategy focuses computational resources in Stage 2 on problems that are most beneficial for learning.
@@
Qwen2.5-72B-Instruct 11.42 $\pm$ 0.24 31.0 14.50 $\pm$ 1.33 40.0 14.63 $\pm$ 1.03 36.6
+ RFT @ 65k (ours) 20.46 $\pm$ 0.42 43.0 22.50 $\pm$ 1.18 45.0 20.98 $\pm$ 1.51 43.9
+ Stage 1 RL @ 65k (ours) 35.74 $\pm$ 0.28 54.6 36.50 $\pm$ 1.59 55.0 31.22 $\pm$ 0.80 53.7
+ Stage 2 RL @ 131k (ours) 39.04 $\pm$ 0.50 58.4 35.00 $\pm$ 1.54 52.5 31.71 $\pm$ 1.31 53.7
DeepSeek-V3-0324 39.56 $\pm$ 0.47 62.2 36.75 $\pm$ 0.92 60.0 31.46 $\pm$ 1.38 58.5
Qwen3-235B no-thinking 25.84 $\pm$ 0.37 54.4 27.25 $\pm$ 1.15 57.5 22.93 $\pm$ 2.16 48.8
Qwen3-32B no-thinking 20.40 $\pm$ 0.34 44.0 21.75 $\pm$ 1.54 50.0 17.56 $\pm$ 1.30 36.6
Llama4 Maverick 15.84 $\pm$ 0.54 47.2 19.00 $\pm$ 1.72 50.0 13.66 $\pm$ 1.79 39.0
4 Training Details and Results

### v2

@@ Abstract
Research on applications of reinforcement learning (RL) to large language models has mostly been focused on single-turn problems, such as mathematical reasoning or single-shot code generation. While these problems can be viewed as token-level multi-turn Markov decision processes (MDPs), this view corresponds to a degenerate case of multi-turn interaction where the environment provides no feedback. This contrasts with many real-world domains, such as software engineering (SWE), which require rich multi-turn interactions with a stateful environment that responds to each action with a non-trivial observation. To bridge this gap, we demonstrate the successful application of RL to this general regime. Our methodology begins with rejection fine-tuning (RFT) using execution feedback to train a policy to follow instructions and formatting effectively, followed by a synchronous RL pipeline using DAPO for iterative improvement. Applying this pipeline to Qwen2.5-72B-Instruct, we increase its Pass@1 on the SWE-bench Verified benchmark from 11% to 39%, substantially improving upon the 20% RFT baseline. On the May and June splits of SWE-rebench, the resulting agent achieves Pass@1 of 35% and 31% respectively, competitive with even larger models such as DeepSeek-V3-0324 or Qwen3-235B-A22B, demonstrating that our methodology offers a practical approach for training capable agents for multi-turn interactive tasks using open-weight models.
@@ 1 Introduction
Large language models (LLMs) are increasingly being deployed within autonomous agents in complex, real-world domains. Software engineering (SWE) is an especially compelling application area, promising substantial economic impact through automation of debugging, code generation, and software maintenance tasks. However, current approaches to developing effective SWE agents rely predominantly on one of three strategies: (i) combining sophisticated scaffolding with proprietary LLMs (AugmentCode, 2025; Refact AI, 2025; Trae, 2025), (ii) leveraging extensive inference-time scaling techniques (Antoniades et al., 2025), or (iii) supervised fine-tuning (SFT) of open-weight models using demonstrations from stronger teacher models (Yang et al., 2025; Zeng et al., 2025; Pan et al., 2025). While these methods have yielded strong initial results, they are often resource-intensive and depend on powerful, proprietary models. This reality underscores the need for methods that can build comparably effective systems from smaller, open-weight models. Reinforcement learning (RL) offers a promising alternative by directly optimizing an agent’s policy through interaction with a responsive environment, potentially achieving stronger performance without reliance on teacher models.
@@
Figure 1: Illustration of task structure differences between bandit-style problems (top, e.g., math) and POMDPs (bottom, e.g., software engineering), defined in Section 3.1. In bandit settings, the agent takes a single action to produce a final solution based on an initial observation. In contrast, POMDPs require a multi-step interaction loop where the agent repeatedly takes actions and interprets new environmental feedback to guide its subsequent decisions.
@@
  • Data scalability and fidelity: Generating high-quality trajectories requires the reproduction of specific repository states in controlled environments, which limits dataset scale. Large-scale datasets such as SWE-smith (Yang et al., 2025) and SWE-rebench (Badertdinov et al., 2025) begin to address this gap; we primarily build on the latter.
@@
In this paper, we address these challenges by developing and validating a complete training pipeline for interactive SWE agents. Our primary contributions are:
  • We design and implement a robust two-phase training methodology that combines rejection fine-tuning (RFT) with a subsequent reinforcement learning stage using DAPO (Yu et al., 2025). This pipeline appears to be effective for training agents in long-context interactive environments.
  • We provide a strong empirical demonstration of our method’s effectiveness. By applying our pipeline to a 72B-parameter open-weight model, we improve its Pass@1 on SWE-bench Verified (Chowdhury et al., 2024) from 11% to 39.0%. Coupled with 35% and 31% Pass@1 scores on the SWE-rebench May/June splits, the agent achieves competitive performance with powerful, often much larger, open-weight models (Table 1).
  • We share key practical insights and analyses from scaling RL to long-context SWE tasks. This includes findings on maintaining training stability and lessons learned that can serve as a blueprint for future work on RL for complex, interactive agents.
2 Related work
Software engineering agents.
Early systems, notably SWE-agent (Yang et al., 2024), demonstrated that LLMs could effectively operate within sandboxed software environments using predefined toolkits (e.g., shell commands, code editors) and be evaluated via automated unit tests (Jimenez et al., 2024). Subsequent frameworks introduced alternative scaffolding and prompting strategies to enhance model interactions, including Agentless (Xia et al., 2024), OpenHands (Wang et al., 2025), and Moatless (Antoniades et al., 2025). Our work closely aligns with the original SWE-agent design, leveraging similar prompting structures and tool configurations, enabling a direct evaluation of reinforcement learning improvements within this established context.
Strategies to improve SWE agents.
Prior efforts to enhance SWE agent performance beyond improving scaffoldings broadly focus on either test-time exploration or model-level improvements through supervised training. Test-time exploration strategies, such as Monte Carlo Tree Search (MCTS) (Antoniades et al., 2025) and guided 1-step lookahead (Zainullina et al., 2025), significantly boost task success by exploring multiple solution trajectories. However, these methods are computationally expensive and sometimes introduce infrastructure complexity due to operations such as environment rollbacks and parallel execution paths.
Alternatively, many efforts have targeted model-level improvements via supervised fine-tuning using expert-curated demonstrations. Prominent examples include SWE-smith (Yang et al., 2025), SWE-Fixer (Xie et al., 2025), and Skywork-SWE (Zeng et al., 2025), all of which achieved success by training open-weight models on extensive demonstration data. In contrast, our method relies exclusively on self-generated interaction data obtained through direct RL training. This approach simplifies data collection, eliminates the need for strong teacher models and opens the way for iterative self-improvement.
Reinforcement learning for coding.
RL has shown notable success in structured reasoning domains like mathematics (Shao et al., 2024; Seed et al., 2025). In the coding domain, early RL applications often focused on single-turn code generation (Dou et al., 2024). In complex SWE tasks, SWE-RL (Wei et al., 2025) is a recent example of applying policy-gradient RL, where the agent achieved strong results. However, its methodology, built upon the Agentless scaffold, frames the problem as a single-turn task where the model generates a complete solution patch in one pass. This approach simplifies the learning problem by sidestepping the complexities of stateful, multi-turn interaction and long-horizon credit assignment.
Our work, alongside other recent advancements, directly addresses these multi-turn challenges. Concurrent to our work, DeepSWE (Luo et al., 2025) successfully scaled critic-free RL training to a 32B-parameter model, while Sky-RL (Cao et al., 2025) introduced an asynchronous RL pipeline for long-context tasks. Our contribution to this emerging area is a complete, multi-stage methodology that successfully applies RL to a larger 72B-parameter model with a 131k token context in a fully interactive, multi-turn setting.
@@ 3.1 Task formulation
We formalize the task of an autonomous SWE agent as a partially observable Markov decision process (POMDP) (Murphy, 2025), defined by the tuple $\langle\mathcal{Z},\mathcal{A},\Omega,T,\mathcal{O},R,\gamma\rangle$:
@@
Our goal is to find a policy $\pi_{\theta}$ that maximizes the expected cumulative reward. The cumulative reward for a trajectory $\tau$ is the sum of rewards over all its steps, $G(\tau)=\sum_{t=0}^{|\tau|-1}R(z_{t},a_{t})$, where $|\tau|$ denotes the number of actions in the trajectory. Given our sparse reward structure where immediate rewards $R(z_{t},a_{t})$ are zero for all non-terminal steps, this simplifies to the terminal reward, which we denote as $R(\tau)$. $R(\tau)$ is $1$ if the final proposed patch passes the validation test suite, and 0 otherwise.
@@ 3.2 From PPO to DAPO
Reinforcement learning has traditionally been dominated by Proximal Policy Optimization (PPO) (Schulman et al., 2017), which employs a learned critic to estimate the advantage of each generated action, shown in Equation 5. While powerful, PPO introduces extra overhead and sensitivity to hyperparameter tuning.
Group-Relative Policy Optimization (GRPO) (Shao et al., 2024) eliminates the need for a learned advantage estimator. Instead, it computes a Monte Carlo estimate of the advantage for each trajectory by normalizing its terminal reward against the mean and standard deviation of rewards from a group of trajectories sampled from the same policy.
@@
where $\sigma_{R}$ is the standard deviation of rewards in the group and $\delta$ is a small constant for numerical stability. This single advantage value $\hat{A}^{(i)}$ is then applied to all tokens within the trajectory $\tau^{(i)}$, shown in Equation 7.
@@
  • Dynamic sampling: Sampled trajectories that carry no learning signal (i.e., $\hat{A}^{(i)}=0$) are filtered out, focusing computation on effective updates.
  • Soft overlong punishment: An incremental penalty is imposed on trajectories that exceed a predefined threshold. The penalty scales linearly within a specified interval and is added to the original test-based reward to discourage excessively long responses.
@@
Our implementation applies the DAPO framework, adapting the reward function to better suit the multi-turn nature of SWE tasks, as detailed in Section 4.3. Complete mathematical formulations for objective functions are provided in Appendix A.
@@ 3.3 Agent scaffolding
Our agent follows the SWE-agent (Yang et al., 2024) implementation with a ReAct-style loop (Yao et al., 2023), interacting with the environment through a predefined set of tools. The entire action–observation history conditions every decision. The agent interacts with an environment using the following tools:
@@
Each SWE task includes a GitHub-style issue with a natural language description, a failing test suite that evaluates final patch correctness, and a sandboxed environment initialized from a repository snapshot. Prompts, available tools and an example of an SWE task can be found in Appendices F-G.
4 Training methodology
This section outlines our carefully curated data and a two-phase training pipeline optimized for multi-turn SWE tasks.
@@ 4.1 Data
  • Deterministic tests: Remove tasks with flaky tests that produce inconsistent results across repeated executions (50 trials), ensuring stable training signals. Non-deterministic test behavior occurs mostly due to external service calls or floating-point inaccuracies.
For evaluation, we use the standard SWE-bench Verified benchmark, a 50-problem random subset (referred to as Verified-50) for faster intermediate checkpoint evaluation, and the monthly splits of SWE-rebench (May-June) which are not included in the training set, ensuring a fair and contamination-free comparison. The full list of Verified-50 problems can be found in Appendix B.
@@ 4.2 Phase 1: rejection fine-tuning (RFT)
We start from the open-weight Qwen2.5-72B-Instruct model. Out of the box it achieves only $\sim$11% Pass@1 on SWE-bench Verified with poor instruction following being the dominant issue, resulting in improperly formatted commands. Examples of such behavior can be found in Appendix E.
To address this problem and to warm up the model for the RL stage, we perform rejection fine-tuning. First, we run the initial checkpoint 10 times on the selected SWE-rebench tasks and keep only trajectories whose patches pass the test suite. This yields a set of 6,548 successful trajectories, on which we run a single epoch of supervised fine-tuning. During this epoch, we mask assistant turns that triggered an environment-formatting error, thereby focusing the loss only on valid actions and improving adherence to the tool structure (see Figure 2). After RFT, the model’s accuracy rises to $\sim$20% (Table 1) and serves as a baseline for comparison with subsequent RL runs. Hyperparameters for the RFT run are listed in Appendix C.
Figure 2: An example trajectory from the agent’s interaction used in RFT. Only green (error-free) assistant turns contribute to training loss.
@@ 4.3 Phase 2: multi-turn RL
The core of our work involves applying RL to thousands of problems in an iterative loop. Each RL iteration includes:
@@
  • Reward computation: Following DAPO formulation, the final reward signal $R_{\text{final}}(\tau)$ combines the binary success reward from test execution with a trajectory length penalty. Unlike the original token-based penalty, we apply a linear penalty for exceeding a predefined number of steps to better reflect the multi-turn nature of SWE tasks. The reward is computed as follows:
@@
    Here, $R(\tau)\in\{0,1\}$ is the terminal reward defined in Section 3.1, $L_{\text{thr}}$ and $T_{\text{max}}$ are hyperparameters representing the penalty threshold and maximum number of turns per trajectory, respectively.
@@
  • Optimization: We update all model parameters using DAPO’s clipped token-level objective.
Our RL training is divided into two sequential stages. The first stage establishes a baseline policy at a 65k context length. The second stage then advances this policy by training on a longer 131k context and doubling the maximum number of agent turns $T_{\text{max}}$, allowing the agent to tackle more complex problems.
Transitioning to this more computationally demanding setting required adjusting key hyperparameters to ensure stable and efficient training. Following practices from training reasoning models with large-scale RL (Rastogi et al., 2025; Parashar et al., 2025; He et al., 2025), we decreased the high clip bound, increased problem difficulty and batch size, and decreased the number of instances sampled per iteration. The problems difficulty is increased by removing instances that have a success rate of $2/3$ over all training iterations. To eliminate instances that are likely unsolvable, we also filter those that have never been solved. The complete set of hyperparameter changes between stages is detailed in Table 3.
@@
Qwen2.5-72B-Inst 11.4 $\pm$ 0.24 31.0 14.5 $\pm$ 1.33 40.0 14.6 $\pm$ 1.03 36.6
+ RFT @ 65k 20.5 $\pm$ 0.42 43.0 22.5 $\pm$ 1.18 45.0 21.0 $\pm$ 1.51 43.9
+ Stage 1 RL @ 65k 35.7 $\pm$ 0.28 54.6 36.5 $\pm$ 1.59 55.0 31.2 $\pm$ 0.80 53.7
+ Stage 2 RL @ 131k 39.0 $\pm$ 0.50 58.4 35.0 $\pm$ 1.54 52.5 31.7 $\pm$ 1.31 53.7
Llama-4 Maverick 15.8 $\pm$ 0.54 47.2 19.0 $\pm$ 1.72 50.0 13.7 $\pm$ 1.79 39.0
Qwen3-32B no-think 20.4 $\pm$ 0.34 44.0 21.8 $\pm$ 1.54 50.0 17.6 $\pm$ 1.30 36.6
gpt-oss-120b 22.8 $\pm$ 0.54 58.4 24.8 $\pm$ 1.26 57.5 19.5 $\pm$ 2.36 48.8
Qwen3-235B no-think 25.8 $\pm$ 0.37 54.4 27.3 $\pm$ 1.15 57.5 22.9 $\pm$ 2.16 48.8
DeepSeek-V3-0324 39.6 $\pm$ 0.47 62.2 36.8 $\pm$ 0.92 60.0 31.5 $\pm$ 1.38 58.5
Qwen3-235b-Inst-2507 46.9 $\pm$ 0.19 69.4 41.3 $\pm$ 1.13 57.5 38.3 $\pm$ 1.93 61.0
5 Results and findings